# LiRA: Linguistic Robust Anchoring for Cross-lingual Large Language Models

### v1

@@ Abstract
As large language models (LLMs) rapidly advance, performance on high-resource languages (e.g., English, Chinese) is nearing saturation, yet remains substantially lower for low-resource languages (e.g., Urdu, Thai) due to limited training data, machine-translation noise, and unstable cross-lingual alignment. We introduce LiRA (Linguistic Robust Anchoring for Large Language Models), a training framework that robustly improves cross-lingual representations under low-resource conditions while jointly strengthening retrieval and reasoning. LiRA comprises two modules: (i) Arca (Anchored Representation Composition Architecture), which anchors low-resource languages to an English semantic space via anchor-based alignment and multi-agent collaborative encoding, preserving geometric stability in a shared embedding space; and (ii) LaSR (Language-coupled Semantic Reasoner), which adds a language-aware lightweight reasoning head with consistency regularization on top of Arca’s multilingual representations, unifying the training objective to enhance cross-lingual understanding, retrieval, and reasoning robustness. We further construct and release a multilingual product retrieval dataset covering five Southeast Asian and two South Asian languages. Experiments across low-resource benchmarks (cross-lingual retrieval, semantic similarity, and reasoning) show consistent gains and robustness under few-shot and noise-amplified settings; ablations validate the contribution of both Arca and LaSR. Code will be released on GitHub and the dataset on Hugging Face.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress across a wide range of natural language understanding and reasoning tasks. However, their performance remains heavily skewed toward high-resource languages such as English and Chinese, while performance on low-resource languages (LRLs) continue to lag far behind. This disparity primarily arises from long-tailed pretraining distributions Conneau et al. (2020) (Figure 1a), insufficient or noisy parallel data Khayrallah & Koehn (2018), and unstable cross-lingual alignment Wu & Dredze (2020). Consequently, the direct application of LLMs to retrieval and reasoning tasks in LRLs frequently results in diminished performance and inconsistent behavior, thereby compromising the potential for achieving global inclusiveness.
@@
Existing approaches to cross-lingual adaptation typically rely on machine translation Artetxe et al. (2023); Shubham (2024) or multilingual Singh et al. (2024); Mei & Zhao (2025). While translation-based pipelines have demonstrated a certain degree of effectiveness, they are inherently susceptible to error propagation and semantic drift, particularly in complex tasks involving multi-step reasoning or nuanced linguistic interpretation. On the other hand, multilingual encoders provide a degree of language-agnostic representation but lack the strong reasoning capability that LLMs demonstrate in English. Although prior work, such as MindMerger Huang et al. (2024b), has made progress in narrowing the performance gap of large language models in low-resource languages, its architectural design remains relatively simplistic and lacks grounding in a rigorous theoretical framework, limiting its generalizability and analytical depth. Similarly, Lusifer Man et al. (2025) proposes a straightforward integration of a multilingual encoder with an LLM-based embedding model, primarily targeting embedding tasks rather than cross-lingual reasoning. Despite their empirical gains, both approaches remain theoretically underdeveloped. As a result, neither approach adequately closes the performance gap for low-resource languages across key tasks such as information retrieval, ranking, and reasoning.
To address these challenges, we introduce LiRA (Linguistic Representation Anchoring), a unified framework that anchors low-resource languages to the English semantic space while preserving LLM-level reasoning ability. Unlike prior approaches, LiRA is grounded in a rigorous theoretical foundation that certifies the completeness and stability of its cross-lingual representations (Figure 1b). LiRA integrates two complementary components: (i) Arca, which explicitly aligns multilingual representations with English through critic–actor interaction and feature anchoring, thus reducing semantic drift; and (ii) LaSR, a lightweight head that fuses multilingual and English embeddings with queue-based objectives, ensuring robust retrieval and reasoning under low-resource conditions. In this way, LiRA not only leverages the strong English capability of LLMs but also transfers it effectively to underrepresented languages.
@@
  • We propose a novel cross-lingual reasoning framework that generalizes the strong English ability of LLMs to mid- and low-resource languages.
@@
  • We release a retrieval dataset covering 5 Southeast Asian and 2 South Asian mid- and low-resource languages, enabling further research in this underexplored area.
@@ 2 Related Work
2.1 Cross-lingual Information Retrieval
Early CLIR systems relied on multilingual PLMs (e.g., mBERT Devlin et al. (2019), XLM-R Conneau et al. (2020)) for alignment; recent work moves toward supervision-light transfer and LLM-embedding adaptation. LUSIFER integrates a multilingual encoder with an English-specialized LLM-based embedding model via a connector to yield strong zero-shot multilingual retrieval Man et al. (2025). CORA unified cross-lingual dense retrieval with target-language generation for many-to-many QA, shaping the CLIR+QA paradigm Asai et al. (2021); later studies improved CLIR with sequential relation modeling and unsupervised multilingual dense retrieval Zhang et al. (2023); Huang et al. (2024a). Beyond passage-level retrieval, CCPR formulates contextualized phrase-level cross-lingual retrieval to mitigate polysemy Li et al. (2024), while XRAG benchmarks end-to-end cross-lingual RAG from retrieval to generation Liu et al. (2025). Domain resources (e.g., CrossMath) and synthetic datasets (e.g., SWIM-IR) further broaden evaluation and training coverage Gore et al. (2024); Thakur et al. (2023). Finally, cross-lingual contamination can inflate reported performance, calling for stricter data governance Shi et al. (2024).
2.2 LLM-based Cross-lingual Reasoning
Recent studies have intensified attention to reasoning tasks in low-resource language settings. MindMerger merges external multilingual understanding into LLMs and trains collaborative use of internal reasoning and external language skills, substantially boosting reasoning in non-English settings Huang et al. (2024b). Process-level improvements—such as Chain-of-Preference Optimization and Chain-of-Code—provide transferable reasoning scaffolds that can be combined with language anchoring/forcing strategies Liu et al. (2024); Zhang et al. (2024). Mechanism-focused studies reveal cross-lingual knowledge barriers and even “cross-lingual collapse” of reasoning traces toward dominant pretraining languages; mitigation includes mixed-language finetuning, reward shaping for language consistency, and representation steering to strengthen non-English token representations Chua et al. (2024); Zhao et al. (2024); Park et al. (2025); Mahmoud et al. (2025). Recent surveys systematize multilingual reasoning evaluations and resources Ghosh et al. (2025).
@@ 3.1 Preliminaries
Setup. Let $\mathcal{X}$ denote a low-resource language (LRL) sentence space and $\mathcal{Y}$ the English sentence space. for any $y^{\prime}\in\mathcal{N}_{\delta}(y)$ we set $\mathbf{z}^{\star}\coloneqq[\,g(x);\,h(y^{\prime})\,]$ (with the same $x$), so that the ratio compares $f_{\mathrm{LLM}}$ under $(y,y^{\prime})$. $\mathcal{Y}_{\mathrm{data}}$ is the finite evaluation corpus on which neighborhoods and empirical quantiles are computed. Given a source sentence $x\in\mathcal{X}$, an LLM-based translator $T:\mathcal{X}\!\to\!\mathcal{Y}$ produces $y=T(x)$, and $y^{\star}\in\mathcal{Y}$ denotes a perfect translation that matches $x$ in semantic distribution (formalized below). We consider two representation paths: an anchor map $g:\mathcal{X}\!\to\!\mathbb{R}^{d}$ that embeds LRL sentences directly into the “English” semantic space, and an English encoder $h:\mathcal{Y}\!\to\!\mathbb{R}^{d}$. We form a concatenated cross-lingual representation $\mathbf{z}\!=\![\,g(x);\,h(y)\,]\in\mathbb{R}^{2d}$, which is consumed by a downstream scorer $f_{\text{LLM}}:\mathbb{R}^{2d}\!\to\!\mathbb{R}$ for retrieval or reasoning. Our goal is to certify the fidelity of $\mathbf{z}$ and provide measurable error bounds.
@@
$\|g(x)-h(T(x))\|_{2}\leq\epsilon_{1},\qquad\epsilon_{1}\geq 0.$
@@
Let $h$ denote the English sentence encoder. We model $h(y)$ as the kernel mean embedding (KME) of the semantic distribution $p(s\mid y)$ into an RKHS $\mathcal{H}$ induced by a positive–definite kernel $k$:
$h(y)\;=\;\mu_{p(s\mid y)}\;\coloneqq\;\mathbb{E}_{s\sim p(s\mid y)}\!\bigl[\varphi(s)\bigr],\qquad\varphi(s)\;\coloneqq\;k(s,\cdot).$
@@
for some constant $C>0$. See  A.1.2 for details.
@@
$L_{h}^{\mathrm{loc}}(y;\delta)\;\coloneqq\;\max_{y^{\prime}\in\mathcal{N}_{\delta}(y)}\frac{\bigl\|f_{\mathrm{LLM}}(\mathbf{z})-f_{\mathrm{LLM}}(\mathbf{z}^{\star})\bigr\|_{2}}{\bigl\|\mathbf{z}-\mathbf{z}^{\star}\bigr\|_{2}},\qquad\mathbf{z}\,=\,[\,g(x)\,;\,h(y)\,].$
Here $\mathcal{N}_{\delta}(y)\!=\!\{\,y^{\prime}\in\mathcal{Y}_{\mathrm{data}}:0<d_{\mathrm{tok}}(y,y^{\prime})\leq\delta\,\}$ denotes the token-edit neighborhood (we use $\delta\!=\!1$ in most experiments). We also report the empirical $q$-quantile $L_{h}^{(q)}(\delta)$, where $q\in(0,1)$ is the quantile level; $L_{h}^{(q)}(\delta)$ is the $q$-th empirical quantile of $L_{h}^{\mathrm{loc}}(y;\delta)$ over $y\in\mathcal{Y}_{\mathrm{data}}$. ; for example, with q=0.95q\!=\!0.95 we observe $L_{h}^{(0.95)}\!\approx\!0.034$ (see A.1.2 for details.).
@@ 3.2 Theorem
Theorem (Representation deviation).
Under Assumptions 12 and Definitions 12, let $\mathbf{z}^{\star}=[\,h(y^{\star});\,h(y^{\star})\,]$ and $\mathbf{z}=[\,g(x);\,h(y)\,]$. Then
@@
Corollary (Downstream stability).
If $f_{\mathrm{LLM}}$ is locally Lipschitz with constant $L^{\mathrm{loc}}(y;\delta)$ as in Definition 2, then
@@
In practice, we report empirical estimates, e.g., $L^{(0.95)}\!\approx\!0.034$ and $C\!\approx\!0.6866$ see  A.1.2 for details. The theorem and its corollary guarantee that, under Assumptions 1–2, the model obtains high-fidelity (robust) representations that effectively support downstream tasks. Full math proof of theorem and corollary details are provided in  A.1.
3.3 WHY CONCATENATE TWO REPRESENTATION PATHS?
$g(x)=s+\eta_{g},\qquad h(y)=s+\eta_{h},$
Define the information gain of concatenation:
$\Delta I\;=\;I\!\bigl(s;\,[g(x),h(y)]\bigr)\;-\;I\!\bigl(s;\,g(x)\bigr).$
$I\!\bigl(s;\,[g,h]\bigr)\;=\;I(s;g)\,+\,I(s;h\mid g)\;=\;I(s;g)+H(s\mid g)-H(s\mid g,h)\;\geq\;I(s;g),$
with equality iff $s\!\perp\!\!\!\perp\!h\mid g$. In cross-lingual settings, translation noise $\eta_{h}$ and anchoring noise $\eta_{g}$ are complementary, hence $I(s;h\mid g)\!>\!0$ and $\Delta I\!>\!0$. Therfore, if $\mathrm{Var}(\eta_{g,k})\!\to\!\infty$ on some dimension $k$ (e.g., severe LRL ambiguity), then
$I(s;g)\;\leq\;\tfrac{1}{2}\log\!\Bigl(1+\tfrac{\sigma_{s,k}^{2}}{\mathrm{Var}(\eta_{g,k})}\Bigr)\!\to 0,$
@@ 4.1 ARCA
Let $\mathcal{X}$ be a low–resource source space and $\mathcal{Y}$ the English space. For any $x\in\mathcal{X}$, a translator $T:\mathcal{X}\!\to\!\mathcal{Y}$ yields $y=T(x)$. We introduce two representation paths: an anchoring map $g:\mathcal{X}\!\to\!\mathbb{R}^{d}$ that lands source sentences directly in an “English semantic” space, and an English encoder $h:\mathcal{Y}\!\to\!\mathbb{R}^{d}$. We concatenate $\mathbf{z}=[g(x);\,h(y)]$ and score it with an LLM head $f_{\text{LLM}}$. ARCA aims to reduce the two terms appearing in our generalization bound (Sec. 3.2): the anchoring error $\epsilon_{1}$ and the translation distortion $\epsilon_{2}$.
ARCA (Figure 2) comprises three modules: (i) a Translation Critic that judges candidates with semantic/emotional/pragmatic scores and (projected) embedding similarity; (ii) an Embedding Critic that anchors feature paths to translation paths via a regression-style penalty; and (iii) an Actor trained with policy gradients that fuses both critics.
Figure 2: An overview of Arca.
4.1.1 Embedding Critic and Feature Anchoring (minimizing $\epsilon_{1}$)
On the query side, the input is constructed with a feature prefix of length $S_{\text{feat}}$ followed by text tokens of length $T$: $\mathbf{H}^{(q)}\in\mathbb{R}^{(S_{\text{feat}}+T)\times D}$. On the document (candidate) side we have $\mathbf{H}^{(d)}\in\mathbb{R}^{T\times D}$. We define the feature-path representation (average over the feature prefix)
$\mathbf{g}_{i}\;=\;\frac{1}{S_{\text{feat}}}\sum_{s=1}^{S_{\text{feat}}}\mathbf{H}^{(q)}_{s,:}\,$, (3)
and the translation-path representation as pooled document states
$\mathbf{h}_{i}\;=\;\mathrm{pool}\!\big(\mathbf{H}^{(d)}\big),$ (4)
where $\mathrm{pool}(\cdot)$ (e.g., last-token/mean/attention) eliminates the tokenizer-induced length mismatch; the pooling dimension is a hyperparameter.
The feature-anchoring loss penalizes the discrepancy between the two paths:
$\mathcal{L}_{\text{anchor}}=\frac{1}{B}\sum_{i=1}^{B}\big\|\mathbf{g}_{i}-\mathbf{h}_{i}\big\|_{2}^{2}.$ (5)
Role for $\epsilon_{1}$. Eq. equation 5 directly contracts the anchoring radius $\epsilon_{1}$ by regressing the feature-path vector towards the translation-path vector.
[…]
$\text{sim}_{k}=\cos(\tilde{e},e_{k}).$

### v2

# Toward Robust Multilingual Adaptation of LLMs for Low-Resource Languages

@@ Abstract
Large language models (LLMs) continue to struggle with low-resource languages, primarily due to limited training data, translation noise, and unstable cross-lingual alignment. To address these challenges, we propose LiRA (Linguistic Robust Anchoring for LLMs)—a plug-and-play framework that requires only lightweight fine-tuning on top of existing pretrained backbones. LiRA jointly optimizes representation stability and cross-lingual semantic consistency by combining two key components: Arca (Anchored Representation Composition Architecture), which aligns low-resource inputs to a shared English semantic space through anchor-based alignment and collaborative encoding; and LaSR (Language-coupled Semantic Reasoner), a lightweight, language-aware head that enforces consistency regularization for unified cross-lingual understanding, retrieval, and reasoning. We theoretically show that under controlled anchoring error and translation-induced bias, LiRA guarantees bounded representation deviation and stable downstream performance under local Lipschitz continuity. To facilitate research, we release a new multilingual product retrieval dataset covering five Southeast Asian and two South Asian languages. Extensive experiments across diverse low-resource benchmarks demonstrate consistent improvements in retrieval, ranking, question answering, and reasoning tasks. Code will be publicly available on GitHub, and the dataset will be hosted on Hugging Face.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have made remarkable progress in natural language understanding and reasoning. Yet these gains are unevenly distributed: performance is concentrated in high-resource languages such as English and Chinese, while low-resource languages (LRLs) continue to lag far behind. This gap is driven by long-tailed pretraining distributions (Haddow et al., 2022) (Figure 1a), limited or noisy parallel data (Ataman et al., 2025), and unstable cross-lingual alignment (Xu et al., 2025). As a result, directly deploying LLMs for retrieval and reasoning in LRLs often leads to degraded accuracy and brittle, inconsistent behavior, limiting the promise of truly inclusive NLP systems.
@@
Existing cross-lingual adaptation methods typically fall into two families: machine translation (MT)-based pipelines (Artetxe et al., 2023; Shubham, 2024) and multilingual approaches (Singh et al., 2024). Although MT-based pipelines can be effective, they are prone to error propagation (Wu et al., 2019) and semantic drift (Beinborn and Choenni, 2020), especially for complex settings that require multi-step reasoning or nuanced interpretation. Multilingual encoders, in contrast, offer more language-agnostic representations but often fail to inherit the strong English-centric reasoning ability exhibited by LLMs (Hu et al., 2020). Recent systems such as MindMerger (Huang et al., 2024) attempt to narrow this gap by integrating the capabilities of LLMs and multilingual models, yet its reliance on parallel translation data may still lead to error propagation and semantic drift. Similarly, Lusifer (Man et al., 2025) connects a multilingual encoder with an LLM-based embedding model to enable cross-lingual transfer; however, its training paradigm lacks information from low-resource languages, which may compromise the performance of the transfer during inference. Moreover, despite empirical gains, these lines of work remain theoretically underdeveloped and do not consistently close the gap on key low-resource tasks such as retrieval, ranking, and reasoning.
Our work is motivated by the observation that cross-lingual adaptation still relies heavily on machine translation pipelines or multilingual encoders. In real-world settings, such as low-resource e-commerce, these solutions are often undermined by translation noise and unstable cross-lingual representations. More importantly, they rarely model how semantic drift and representation-mapping errors propagate through the system, making robustness difficult to analyze and even harder to improve systematically. To address these challenges, we introduce LiRA (Linguistic Representation Anchoring), a plug-and-play framework that anchors low-resource languages to an English semantic space while preserving LLM-level reasoning (Figure 1b). LiRA can be attached to a variety of pretrained backbones and requires only lightweight fine-tuning to improve cross-lingual representations. Unlike prior approaches, LiRA is grounded in a rigorous theoretical framework that provides formal guarantees on the completeness and stability of the learned representations. LiRA integrates two complementary components: (i) Arca, which reduces semantic drift by aligning multilingual representations with English via critic–actor interaction and feature anchoring; and (ii) LaSR, a lightweight head that fuses multilingual and English embeddings with queue-based objectives to support robust retrieval and reasoning in low-resource settings. In this way, LiRA leverages strong English capabilities and transfers them effectively to underrepresented languages.
@@
  • We propose LiRA, a plug-and-play cross-lingual framework that transfers LLMs’ strong English capability to mid- and low-resource languages.
@@
  • We release a product retrieval dataset covering 5 Southeast Asian and 2 South Asian mid- and low-resource languages, enabling further research in this underexplored area.
@@ 2 Related Work
Cross-lingual Information Retrieval Early CLIR systems relied on multilingual PLMs (e.g., mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020)) for alignment; recent work moves toward supervision-light transfer and LLM-embedding adaptation. LUSIFER integrates a multilingual encoder with an English-specialized LLM-based embedding model via a connector to yield strong zero-shot multilingual retrieval (Man et al., 2025). However, the zero-shot training paradigm limits further performance gains. Beyond passage-level retrieval, CCPR formulates contextualized phrase-level cross-lingual retrieval to mitigate polysemy (Li and others, 2024), while XRAG benchmarks end-to-end cross-lingual RAG from retrieval to generation (Liu et al., 2025). Domain resources (e.g., CrossMath) and synthetic datasets (e.g., SWIM-IR) further broaden evaluation and training coverage (Gore and others, 2024; Thakur et al., 2024). Existing methods often rely on heuristic designs without a unified theoretical framework, which may restrict their generality and theoretical interpretability. By contrast, our approach is grounded in a principled framework that unifies multilingual latent embedding spaces, mitigates semantic drift, and supports plug-and-play modular training. Furthermore, cross-lingual contamination can inflate reported performance (Yao et al., 2024), prompting us to propose new real-world data for fair comparison.
LLM-based Cross-lingual Reasoning Recent studies (Cueva et al., 2024) have intensified attention to reasoning tasks in low-resource language settings (Kim et al., 2023). MindMerger merges external multilingual understanding into LLMs and trains collaborative use of internal reasoning and external language skills, substantially boosting reasoning in non-English settings (Huang et al., 2024). However, MindMerger relies heavily on parallel translation corpora, which may introduce semantic drift due to translation noise. Process-level improvements—such as Chain-of-Preference Optimization and Chain-of-Code—provide transferable reasoning scaffolds that can be combined with language anchoring strategies (Liu and others, 2024; Zhang and others, 2024). Mechanism-focused studies reveal cross-lingual knowledge barriers and even “cross-lingual collapse” of reasoning traces toward dominant pretraining languages; mitigation includes mixed-language finetuning, reward shaping for language consistency, and representation steering to strengthen non-English token representations (Chua and others, 2024; Zhao et al., 2024; Park et al., 2025; Mahmoud et al., 2025). Recent surveys systematize multilingual reasoning evaluations and resources (Ghosh et al., 2025). These methods typically rely on a single embedding-space alignment primarily designed for reasoning tasks, and their performance may not transfer well to broader settings such as ranking and retrieval. In contrast, our proposed Arca minimizes translation-induced semantic drift, provides better interpretability, and can be extended to support multiple tasks—including reasoning, ranking, and retrieval—thereby improving overall generalization.
@@ 3.1 Preliminaries
Setup. We study cross-lingual retrieval/reasoning under low-resource inputs. Let $\mathcal{X}$ be the low-resource language (LRL) sentence space and $\mathcal{Y}$ the English sentence space. For each $x\in\mathcal{X}$, an LLM-based translator $T:\mathcal{X}\to\mathcal{Y}$ produces an observed translation $y=T(x)$, and we introduce an (unobserved) ideal translation $y^{\star}\in\mathcal{Y}$ that is semantically equivalent to $x$. We consider two representation paths into a shared $d$-dimensional space: a trained multilingual encoder $g:\mathcal{X}\to\mathbb{R}^{d}$ that embeds $x$ directly into an English-aligned semantic space, and an English encoder $h:\mathcal{Y}\to\mathbb{R}^{d}$ that embeds English sentences. Our system forms the concatenated representation $\mathbf{z}(x)\coloneqq[\,g(x);\;h(y)\,]\in\mathbb{R}^{2d}$, which is then fed into a downstream $f$ (e.g., for retrieval, ranking or reasoning). For analysis, we define the ideal reference representation $\mathbf{z}^{\star}(x)\coloneqq[\,h(y^{\star});\;h(y^{\star})\,]\in\mathbb{R}^{2d}$, corresponding to the target behavior where both paths agree on the same English semantics; our goal is to bound the representation deviation $\|\mathbf{z}(x)-\mathbf{z}^{\star}(x)\|_{2}$ and the induced deviation $f(\mathbf{z}(x))-f(\mathbf{z}^{\star}(x))||_{2}$. See A.3 for theoretic explanation for why we concanate two representation paths.
@@
$\|g(x)-h(y^{*})\|_{2}\leq\epsilon_{1},\qquad\epsilon_{1}\geq 0.$
@@
We model $h(y)$ as the kernel mean embedding (KME) of the semantic distribution $p(s\mid y)$ into an RKHS $\mathcal{H}$ induced by a positive–definite kernel $k$:
$h(y)\;=\;\mu_{p(s\mid y)}\;\coloneqq\;\mathbb{E}_{s\sim p(s\mid y)}\!\bigl[\varphi(s)\bigr],\quad\varphi(s)\;\coloneqq\;k(s,\cdot).$
@@
for some constant $C>0$. See Appendix  A.2 for details.
@@
$L^{\mathrm{loc}}(y;\delta)\;\coloneqq\;\max_{y^{\prime}\in\mathcal{N}_{\delta}(y)}\frac{\bigl\|f_{\mathrm{LLM}}(\mathbf{z})-f_{\mathrm{LLM}}(\mathbf{z}^{\star})\bigr\|_{2}}{\bigl\|\mathbf{z}-\mathbf{z}^{\star}\bigr\|_{2}}.$
Here $\mathbf{z}\,=\,[\,g(x)\,;\,h(y)\,],\quad\mathcal{N}_{\delta}(y)\!=\!\{\,y^{\prime}\in\mathcal{Y}_{\mathrm{data}}:0<d_{\mathrm{tok}}(y,y^{\prime})\leq\delta\,\}$ denotes the token-edit neighborhood (we use $\delta\!=\!1$ in most experiments). We also report the empirical $q$-quantile $L^{(q)}(\delta)$, where $q\in(0,1)$ is the quantile level; $L^{(q)}(\delta)$ is the $q$-th empirical quantile of $L^{\mathrm{loc}}(y;\delta)$ over $y\in\mathcal{Y}_{\mathrm{data}}$. ; for example, with q=0.95q\!=\!0.95 we observe $L^{(0.95)}\!\approx\!0.034$. See Appendix A.2 for details.
@@ 3.2 Theorem
Theorem (Representation deviation) Under Assumptions 12 and Definitions 12, let $\mathbf{z}^{\star}=[\,h(y^{\star});\,h(y^{\star})\,]$ and $\mathbf{z}=[\,g(x);\,h(y)\,]$. Then
@@
Corollary (Downstream stability) For $f_{\mathrm{LLM}}$ that is locally Lipschitz with constant $L^{\mathrm{loc}}(y;\delta)$ as in Definition 2, we have
@@
The theorem and its corollary imply that, by minimizing $\epsilon_{1}$ and $\epsilon_{2}$, the model obtains high-fidelity and robust representations that effectively support downstream tasks. Full proofs of the theorem and corollary are provided in Appendix  A.1.
3.3 Theory assumptions and practical validity.
Our theoretical analysis does not require “ideal translations” or “ideal semantic anchoring” to hold in practice. Instead, it only models two ubiquitous sources of error in cross-lingual alignment: (i) representation mapping error between an ideal target vector $z^{\ast}$ and the observed representation $z$, and (ii) translation-induced noise. Here, $z^{\ast}$ is introduced purely as a mathematical reference for zz—not as a strict real-world prerequisite—and the assumptions are used to derive objectives that explicitly target semantic drift and vector-level misalignment. Empirically, we observe that the gap between $z$ and $z^{\ast}$ (measured by similarity) decreases as training proceeds, and the loss consistently drops across different Arca training tasks, supporting the practical relevance of our theoretical formulation. See D.2 for empirically measurement during training process.
Figure 2: An overview of Arca.
Figure 3: An overview of LaSR.
@@ 4.1 Arca
Let $\mathcal{X}$ be a low–resource source space and $\mathcal{Y}$ the English space. For any $x\in\mathcal{X}$, a translator $T:\mathcal{X}\!\to\!\mathcal{Y}$ yields $y=T(x)$. We introduce two representation paths: an anchoring map $g:\mathcal{X}\!\to\!\mathbb{R}^{d}$ that lands source sentences directly in an “English semantic” space, and an English encoder $h:\mathcal{Y}\!\to\!\mathbb{R}^{d}$. We concatenate $\mathbf{z}=[g(x);\,h(y)]$ and score it with an LLM critic. Arca aims to reduce the two terms appearing in our generalization bound (Sec. 3.2): the anchoring error $\epsilon_{1}$ and the translation distortion $\epsilon_{2}$. Arca (Figure 2) comprises three modules: (i) a Translation Critic that judges candidates with semantic/emotional/pragmatic scores; (ii) an Embedding Critic that anchors feature paths to translation paths via a regression-style penalty; and (iii) an Actor trained with policy gradients that fuses both critics.
4.1.1 Feature Anchoring (minimizing $\epsilon_{1}$)
Given $x\!\in\!\mathcal{X}$ with observed translation $y=T(x)\!\in\!\mathcal{Y}$, we align the multilingual path $g(x)$ and the English path $h(y)$ by resolving tokenizer-length and dimensional mismatches via temporal pooling and a shared Adaptor. Let $g_{\text{tok}}(x)\!\in\!\mathbb{R}^{L_{x}\times d_{g}}$ and $h_{\text{tok}}(y)\!\in\!\mathbb{R}^{L_{y}\times d_{h}}$ denote the token-level streams before sentence pooling. We apply a temporal pooling operator with $S_{\text{feat}}$ bins (a hyperparameter) to obtain fixed-length sequences:
$G(x)$ $=P_{S_{\text{feat}}}\!\big(g_{\text{tok}}(x)\big)\in\mathbb{R}^{S_{\text{feat}}\times d_{g}},$ (3)
$H(y)$ $=P_{S_{\text{feat}}}\!\big(h_{\text{tok}}(y)\big)\in\mathbb{R}^{S_{\text{feat}}\times d_{h}}.$
A single shared Adaptor $A(\cdot)$ maps either side into a common $d$-dimensional space (with internal projections when $d_{g}\!\neq\!d_{h}$). We then pool along the temporal axis to form sentence-level vectors:
$E_{lr}$ $=\mathrm{pool}\!\big(A(G(x))\big)\in\mathbb{R}^{d},$ (4)
$E_{en}$ $=\mathrm{pool}\!\big(A(H(y))\big)\in\mathbb{R}^{d}.$
The feature-anchoring objective contracts the path discrepancy by cosine alignment:
$\mathcal{L}_{\text{anchor}}\;=\;1-\cos\!\big(E_{lr},E_{en}\big).$ (5)
Minimizing Eq. (5) reduces the anchoring radius $\epsilon_{1}$ after length normalization (Eq. (3)) and feature alignment (Eq. (4)).